# REACT: Real-time Efficient Attribute Clustering and Transfer for Updatable 3D Scene Graph

### v1

@@ III Problem Formulation
Let us define an identity function $\mathcal{I}(I^{i}_{t},I^{j}_{t+1})\in\{0,1\}$, such that $\mathcal{I}(I^{i}_{t},I^{j}_{t+1})=1$ when $I^{i}_{t}$ re-appears at $t+1$ as $I^{j}_{t+1}$, and $\mathcal{I}(I^{i}_{t},I^{j}_{t+1})=0$ otherwise. In other words, $I^{i}_{t}$ and $I^{j}_{t+1}$ represent the exact same object recorded at different time. Then, the relationship between times $t$ and $t+1$ is defined in terms of sets $M^{t}_{t+1}$, $A^{t}_{t+1}$, and $N^{t}_{t+1}$. The Matched set $M^{t}_{t+1}$ includes object instances that remain in the scene, but might have moved; formally $M^{t}_{t+1}=\{(I_{t}^{i},I_{t+1}^{j})\mid I_{t}^{i}\in\mathcal{O}_{t}\wedge I_{t+1}^{j}\in\mathcal{O}_{t+1}\wedge\mathcal{I}(I_{t}^{i},I_{t+1}^{j})=1\}$. The Absent set $A^{t}_{t+1}\subseteq\mathcal{O}_{t}$ includes those that were in $\mathcal{O}_{t}$ but do not appear at time $t+1$; and the New set $N^{t}_{t+1}\subseteq\mathcal{O}_{t+1}$ those that only appear at time $t+1$.
The objective is to identify the set of instance associations $\{M^{t}_{t+1},A^{t}_{t+1},N^{t}_{t+1}\}$ between the object sets $\mathcal{O}_{t}$ and $\mathcal{O}_{t+1}$ and update our map such that all instances in $M^{t}_{t+1}$ have their position updated, all instances in $A^{t}_{t+1}$ are removed, and all instances in $N^{t}_{t+1}$ are added.
@@ V-B Learning for Attribute Clustering and Transfer
As demonstrated in Fig. LABEL:fig:cover, the dataset for training the model is formed after the robot’s first visit to an environment. Using the framework described in Sec. V-A, we construct a 3DSG with object nodes containing raw RGB images of objects from different viewpoints. As our methodology focuses exclusively on the appearance of objects; thus, we mask out background elements in all comparison images. We interactively group objects with similar appearances into instance clusters, create training and validation datasets, and train the model with them. Empirically, we train the model for 30 epochs using Adam as the optimizer, with an initial learning rate of 0.001. Since the robot may observe some objects only briefly during its first scan, we augment the training dataset by horizontal flipping followed by random rotations between $-10^{\circ}$ and $10^{\circ}$ with uniform probability.
@@ VII Discussions
While REACT represents a promising step towards adaptability for mobile robots in changing environments, it still has some limitations. A common failure mode for REACT stems from errors in scene graph construction, such as splitting a single object into multiple object nodes, merging nearby objects into one, or failing to register some objects. Accurately constructing object nodes in a scene graph is particularly challenging in environments with many adjacent objects, especially when there is noisy instance segmentation and unreliable sensor depth maps.

### v2

@@ III Problem Formulation
Let us define an identity function $\mathcal{I}(I^{i}_{t},I^{j}_{t+1})\in\{0,1\}$, such that $\mathcal{I}(I^{i}_{t},I^{j}_{t+1})=1$ when $I^{i}_{t}$ re-appears at $t+1$ as $I^{j}_{t+1}$, and $\mathcal{I}(I^{i}_{t},I^{j}_{t+1})=0$ otherwise. In other words, $I^{i}_{t}$ and $I^{j}_{t+1}$ represent the exact same object recorded at different time. Then, the relationship between times $t$ and $t+1$ is defined in terms of sets $M_{t:t+1}$, $A_{t:t+1}$, and $N_{t:t+1}$. The Matched set $M_{t:t+1}$ includes object instances that remain in the scene, but might have moved; formally $M_{t:t+1}=\{(I_{t}^{i},I_{t+1}^{j})\mid I_{t}^{i}\in\mathcal{O}_{t}\wedge I_{t+1}^{j}\in\mathcal{O}_{t+1}\wedge\mathcal{I}(I_{t}^{i},I_{t+1}^{j})=1\}$. The Absent set $A_{t:t+1}\subseteq\mathcal{O}_{t}$ includes those that were in $\mathcal{O}_{t}$ but do not appear at time $t+1$; and the New set $N_{t:t+1}\subseteq\mathcal{O}_{t+1}$ those that only appear at time $t+1$.
The objective is to identify the set of instance associations $\{M_{t:t+1},A_{t:t+1},N_{t:t+1}\}$ between the object sets $\mathcal{O}_{t}$ and $\mathcal{O}_{t+1}$ and update our map such that all instances in $M_{t:t+1}$ have their position updated, all instances in $A_{t:t+1}$ are removed, and all instances in $N_{t:t+1}$ are added.
@@ V-B Learning for Attribute Clustering and Transfer
As demonstrated in Fig. LABEL:fig:cover, the dataset for training the model is formed after the robot’s first visit to an environment. Using the framework described in Sec. V-A, we construct a 3DSG with object nodes containing raw RGB images of objects from different viewpoints. As our methodology focuses exclusively on the appearance of objects; thus, we mask out background elements in all comparison images. We interactively group objects with similar appearances into instance clusters, create training and validation datasets, and train the model with them. Empirically, we train the model for 30 epochs using Adam as the optimizer, with an initial learning rate of 0.001. Since the robot may observe some objects only briefly during its first scan, we augment the training dataset by horizontal flipping followed by random rotations between $10^{\circ}$ and $10^{\circ}$ with uniform probability.
@@ VII Discussions
While REACT represents a promising step towards adaptability for mobile robots in changing environments, it still has some limitations. A common failure mode for REACT stems from errors in scene graph construction, such as splitting a single object into multiple object nodes, merging nearby objects into one, or failing to register some objects. Accurately constructing object nodes in a scene graph is particularly challenging in environments with many adjacent objects, especially when there is noisy instance segmentation and unreliable sensor depth maps. Potential areas for improvement include, but are not limited to, utilizing more accurate segmentation models, applying multi-resolution object representation to allow for a more fine-grained representation of smaller objects and avoid faulty object merging [12], and sensor fusion techniques to achieve more accurate point clouds.